### v1

# Scalable Out-of-Distribution Robustness in the Presence of Unobserved Confounders

Abstract
We consider the task of out-of-distribution (OOD) generalization, where the distribution shift is due to an unobserved confounder ($Z$) affecting both the covariates ($X$) and the labels ($Y$). In this setting, traditional assumptions of covariate and label shift are unsuitable due to the confounding, which introduces heterogeneity in the predictor, i.e., $\hat{Y}=f_{Z}(X)$. OOD generalization differs from traditional domain adaptation by not assuming access to the covariate distribution ($X^{\text{te}}$) of the test samples during training. These conditions create a challenging scenario for OOD robustness: (a) $Z^{\text{tr}}$ is an unobserved confounder during training, (b) $P^{\text{te}}(Z)\neq P^{\text{tr}}(Z)$, (c) $X^{\text{te}}$ is unavailable during training, and (d) the posterior predictive distribution depends on $P^{\text{te}}(Z)$, i.e., $\hat{Y}=E_{P^{\text{te}}(Z)}[f_{Z}(X)]$. In general, accurate predictions are unattainable in this scenario, and existing literature has proposed complex predictors based on identifiability assumptions that require multiple additional variables. Our work investigates a set of identifiability assumptions that tremendously simplify the predictor, whose resulting elegant simplicity outperforms existing approaches.
1 Introduction
In this paper, we explore a class of out-of-distribution (OOD) tasks that involve an unobserved confounder $Z$, introducing data heterogeneity and confounding the relationship between covariates $X$ and the label $Y$. This heterogeneity leads the predictor $\hat{Y}=f_{Z}(X)$ to vary with $Z$, resulting in a shift in the distribution $P(X,Y)$ when $P(Z)$ changes. Such shifts often cause poor generalization for models trained using empirical risk minimization (Quiñonero-Candela et al., 2022).
Addressing OOD tasks in the presence of latent confounders is particularly challenging because $Z$ is unobserved during both training and testing, its distribution shifts ($P^{\text{tr}}(Z)\neq P^{\text{te}}(Z)$), and the posterior predictive distribution depends on $P^{\text{te}}(Z)$ (i.e., $\hat{Y}=E_{Z^{\text{te}}}[f_{Z}(X)]$).
Existing approaches typically focus on distribution shifts in $X$ or $Y$, assuming $Z$ affects only one of them (Shimodaira, 2000; Lipton et al., 2018; Arjovsky et al., 2019; Sagawa et al., 2020; Krueger et al., 2021; Liu et al., 2021b). Recent studies (Alabdulmohsin et al., 2023; Tsai et al., 2024) handle cases where the unobserved confounder $Z$ shifts. These methods rely on additional variables such as proxy and concept variables, where the concept variable $C$ serves as a concept bottleneck, mediating the dependence between $X$ and $Y$ in subgroups indexed by $Z$. However, in many practical cases, a concept variable might not exist or be observable. Furthermore, these methods often rely on complex and hard-to-train techniques, such as generative models or kernel methods  (Lu et al., 2014; Lucas et al., 2019; Paulus et al., 2020).
Motivated by these challenges, we propose a simpler and scalable approach which requires only one additional variable. Inspired by Box’s principle that “all models are wrong, but some are useful,” we introduce a set of identifiability assumptions that simplify the estimator for learning the latent distribution. Unlike prior work, we do not require a concept variable and demonstrate that identifiability is achievable with only a proxy variable or multiple data sources. Based on our identifiability theory, we introduce a novel regularizer, which is crucial for ensuring the convergence of our model, as shown in Section 6.
Contributions. We propose a two-stage method for generalizing to a shift in the unobserved latent confounder using proxies. First, we learn the posterior distribution of the latent confounder given the covariates ($P^{\text{tr}}(Z|X)$) by training an encoder-decoder network that takes $X$ as input and outputs a distribution over the proxy variable $P(S|X)$. We prove that this distribution is identifiable under some assumptions. Using the learned latent distribution, we then train a mixture-of-experts model (MoE), with one expert per confounder assignment $f_{Z}(X)$, trained with a specially-designed regularizer (Theorem 2). During inference, we learn $P^{\text{te}}(Z|X)$ by addressing the shift in $Z$ as a label shift problem and re-weighting $P^{\text{tr}}(Z|X)$ accordingly. We then apply the mixture model to the re-weighted distribution. Our method relies only on either having access to a proxy variable or multiple unlabeled data sources. Despite its simplicity and scalability, our approach outperforms existing methods.
2 Related Work
Invariant OOD methods
Invariant OOD methods focus on learning representations that remain stable across training and test environments. Techniques such as IRM  (Arjovsky et al., 2019), GroupDRO  (Sagawa et al., 2020), HRM  (Liu et al., 2021b), and V-REx  (Krueger et al., 2021) leverage environment labels to train an invariant predictor. Other invariant approaches do not require environment labels  (Nam et al., 2020; Creager et al., 2021; Liu et al., 2021a; Nam et al., 2022). Liu et al. (2024b) relax the strict invariance assumption but still focus on a single predictor for all environments. Additionally, Kaur et al. (2022) extend invariant prediction to multiple shifts. Nevertheless, empirical studies have demonstrated that such methods may underperform compared to empirical risk minimization  (Rosenfeld et al., 2020; Gulrajani & Lopez-Paz, 2020), and theoretically, the invariant predictor assumption does not hold in many scenarios  (Schrouff et al., 2022; Lin et al., 2022).
Proxy Methods
Proxy-based approaches have recently been employed to address sub-population shifts  (Alabdulmohsin et al., 2023; Tsai et al., 2024), which assume access to proxy and concept variables to make accurate predictions under environmental shifts. Tsai et al. (2024) eliminate the need for concept variables if multiple labeled domains with proxy variables are accessible. However, these methods assume that proxy variables are observable in the test data and that test data are available during training. Proxy methods have also been explored in causal inference  (Miao et al., 2018; Tchetgen et al., 2020; Xu & Gretton, 2023), where Miao et al. (2018) and Tchetgen et al. (2020) establish identifiability when multiple proxies for the environment are observed. Xu & Gretton (2023) assume the label $Y$ is a deterministic function of $X$ and $Z$. Our approach differs by relaxing several assumptions: (1) we require only one additional variable - a proxy variable or access to multiple training domains; (2) we do not require access to the test distribution during training; (3) proxy variables are not required in the test set; (4) we allow both causal and anti-causal relationships between features and the label; and (5) we offer a simpler algorithm that avoids generative modeling or kernel methods.
Domain Adaptation
Domain adaptation methods typically assume access to the test distribution during training. Many of these methods seek to learn an invariant feature space by assuming a stable predictor  (Sun & Saenko, 2016; Ganin et al., 2016). Garg et al. (2023) consider for relaxed-label shift, where $P(y)$ shifts and $P(x|y)$ changes in a constrained manner. Test-time adaptation methods have recently been developed to modify models during inference to match the test distribution  (Zhang et al., 2021; Wang et al., 2020; Sun et al., 2020; Niu et al., 2023). For instance, Zhang et al. (2021) adopt a meta-learning strategy, while Wang et al. (2020) and Sun et al. (2020) retrain the model to minimize entropy and use self-supervision at test time, respectively. Recently, test-time adaptation methods addressing spurious correlations have been proposed  (Sun et al., 2023; Eastwood et al., 2024). However, these methods do not address confounder shifts  (Alabdulmohsin et al., 2023).
Causality in OOD and Domain Adaptation
Causal reasoning has been applied to OOD generalization and domain adaptation problems  (Schölkopf et al., 2012; 2021; Christiansen et al., 2021; Sun et al., 2021; Lu et al., 2021; Zhang et al., 2015; Mahajan et al., 2021). Causal models help elucidate distribution shifts and identify invariant relationships  (Schölkopf et al., 2012; Christiansen et al., 2021). Sun et al. (2021) assume that latent space can be disentangled into spurious and causal features, learning the causal latent space from multiple labeled sources to produce a stable predictor. Wang et al. (2022) re-weight data to remove confounding effects between labels and covariates, assume $Y\rightarrow X$ and require several training domains for identifiability. However, unobserved confounders are not considered in these works.
3 Notation and Problem Formulation
$Z$$S$$X$$Y$
(a) Causal DAG for proxy task
$Z$$U$$X$$Y$
(b) Causal DAG for multi-source task
Figure 1: Causal diagrams. The shaded circle denotes unobserved variable and the solid circle denotes observed variable. $X$ is the covariate, $Y$ is the label, $Z$ is the confounder, $S$ is the proxy for the confounder (in the proxy task), and $U$ is the dataset id for the multi-source task.
In this section, we introduce our tasks and formally define our notation. Let $X\in{\mathbb{X}}\subseteq\mathbb{R}^{n_{x}}$ be the input variable (e.g., an image, a graph, a set), $Y\in{\mathbb{Y}}$ be the target variable, and $Z\in[n_{z}]$ with $n_{z}\geq 2$ (where $[n_{z}]:=\{1,\ldots,n_{z}\}$) denote the latent confounder, such as socioeconomic status, sociocultural traits, local economic conditions, and other contextual factors. Let $P^{tr}$ and $P^{te}$ denote train and test distributions respectively. We use $P$ for distributions that are invariant across train and test. Under these settings, we consider two tasks, that we give arbitrary names for clarity.
Proxy OOD Task
This task is described by the causal DAG shown in Figure 1(a). We have an additional variable $S$ which is a proxy for the latent confounder (Pearl, 2010; Kuroki & Pearl, 2014). Kuroki & Pearl (2014) describe a setting where participation in a Social Services Program $X$ affects children’s cognitive ability $Y$. The latent socio-economic status $Z$ acts as a confounder. Father’s occupation and family income $S$ is used as a proxy variable for $Z$. The goal is to train a model that generalizes to test data where the distribution of the latent confounder has shifted (see the causal DAG in Figure 1(a)) using these proxies.
Definition 1 (Proxy OOD).
Assume the data generation process for the input variable $X$, the target label $Y$, the latent confounder variable $Z$, and proxy $S$ follows the causal DAG in Figure 1(a). Let the training data $D^{\text{tr}}=\{(x^{i},y^{i},s^{i})\}_{i=1}^{N^{\text{tr}}}\sim P^{\text{tr}}(X,Y,S)$ be collected with respect to the training confounder distribution $P^{\text{tr}}(Z)$, i.e., $P^{\text{tr}}(X,Y,S)=\sum_{Z\in[n_{z}]}P(X,Y,S\mid Z)P^{\text{tr}}(Z)$. In the OOD test data, the confounder distribution shifts, i.e., $P^{\text{tr}}(Z)\neq P^{\text{te}}(Z)$, while the causal mechanism between $X$, $Y$, $Z$, and $S$ remains invariant. The OOD extrapolation goal is to learn a predictor that can accurately predict the missing target labels in a test dataset $D^{\text{te}}=\{x^{i}\}_{i=1}^{N^{\text{te}}}\sim P^{\text{te}}(X)$ with a new test confounder distribution $P^{\text{te}}(Z)$.
We further assume the following conditions:
  1. 1.
    Discrete Proxy: $S\in[n_{s}]$. If $S$ is continuous, we can discretize it.
  2. 2.
    Overlapping confounder support: $\text{supp}(P^{\text{te}}(Z))\subseteq\text{supp}(P^{\text{tr}}(Z))$.
Multi-source OOD Task
This task does not assume any observed proxies but multiple training datasets from different sources or domains with varying distributions of the confounder as shown by the causal DAG in Figure 1(b). We require only one domain to be labeled. The task is to train a model that generalizes to unseen domains with the domain causing a shift in the distribution of the latent confounder $Z$. For example, consider data from various health centers to predict healthcare outcomes such as disease prevalence $Y$ based on attributes like age, gender and diet $X$. Socio-economic factors (latent confounder $Z$) confound the relationship between $X$ and $Y$. Suppose only the data from one health center is labeled due to resource constraints. The goal is to use labeled data from one health center and unlabeled data from several health centers to train a model that can generalize to predict outcomes in an unseen health center.
Definition 2 (Multi-source OOD).
Assume the data generation process for the input variable $X$, the target label $Y$, the latent confounder $Z$, and dataset identifier $U$ follows the causal DAG in Figure 1(b). Let $P_{k}(X,Y)$ denote the distribution for the $k^{\text{th}}$ dataset, i.e., $P_{k}(X,Y)=\sum_{Z\in[n_{z}]}P(X,Y\mid Z)P(Z\mid U=k)$. We assume we have access to $P_{1}(X,Y)$ and $P_{k}(X)$ for $2\leq k\leq n_{u}$. In the OOD test data, we assume we see a new distribution, i.e., $P^{\text{te}}=P_{n_{u}+1}(X)$, while the causal mechanism between $X$, $Y$, and $Z$ remains invariant. The OOD extrapolation goal is to learn a predictor that can accurately predict the missing target labels in a test dataset $D^{\text{te}}=\{x^{i}\}_{i=1}^{N^{\text{te}}}\sim P^{\text{te}}(X)$.
We further assume the following conditions:
  1. 1.
    Overlapping confounder support: $\text{supp}\big{(}P(Z\mid U=n_{u}+1)\big{)}\subseteq\text{supp}\big{(}P(Z\mid U=1)\big{)}$.
  2. 2.
    Distinct domains: $P(Z\mid U=k)\neq P(Z\mid U=l)$.
4 Identifiability
In general, latent confounder distributions with proxy cannot be identified without any assumptions (Pearl, 2010). Our method focuses on assumptions designed to improve practical applicability. Below, we present sufficient assumptions, justify their relevance, and contrast them with assumptions found in existing literature. While we define the assumptions for the Proxy DAG (variable $S$ in Figure 1(a)), the same assumptions apply to the Multi-source OOD task (variable $U$ in Figure 1(b)). Since the values of $Z$ can be arbitrarily permuted, we establish identifiability of the latent distribution upto a permutation of the values.
Assumption 1 (Structural Compatibility).
The causal DAGs in Figure 1 are Markov-compatible with the train and test distributions (Spirtes et al., 2000).
This assumption is commonly used while establishing identifiability results. In our case, this assumption implies the following conditional independencies: (1) for the Proxy task $(S\perp X\mid Z)$, and (2) for the Multi-source task $(U\perp X\mid Z)$.
Assumption 2 (Linearly Independent Distributions).
Let $M$ denote the matrix of conditional probabilities $P(S\mid Z)$, where each entry $M_{zs}$ is given by $P(S=s\mid Z=z)$. We assume that $M$ has full rank, specifically $\text{rank}(M)=n_{z}$.
Previous works use similar full-rank assumptions to establish identifiability  (Alabdulmohsin et al., 2023; Tsai et al., 2024). This assumption ensures the identifiability of the latent confounder $Z$. Without it, degenerate cases where each data point forms its own cluster (i.e., $\forall i,Z_{i}=i$) would prevent identification.
Assumption 3 (Sufficient Support).
We assume that the cardinality of the support of the proxy variables $S$ (denoted $n_{s}$) is at least as large as the support of the latent confounder $Z$ (denoted $n_{z}$), i.e., $n_{s}\geq n_{z}$  (Alabdulmohsin et al., 2023).
This assumption guarantees that the proxy variables contain enough information to distinguish between different confounder assignments (up to permutations).
Assumption 4 (Weak Overlap).
There exists $\eta\in\left(\frac{1}{2},1\right]$ such that, for each $i\in\{1,2,\dots,n_{z}\}$,
$\max_{x}P(Z=i\mid X=x)\geq\eta.$
Intuitively, this assumption suggests that for each latent variable assignment, $Z=i$, there is a specific region in the data where it is dominant. For some observation $X=x$, the probability of $Z=i$ given $X=x$ is greater than $\frac{1}{2}$. This prevents excessive overlap between latent variables, ensuring they can be more easily distinguished based on the observed data.
Next, we show that under these assumptions, the latent distribution is approximately identifiable, with the identifiability error being a function of $\eta$. As $\eta$ increases, indicating less overlap, the error decreases. For large values of $\eta$ (i.e., $\eta\to 1$), the latent distribution becomes increasingly identifiable, and the error approaches zero.
Theorem 1 (Approximate Identifiability of Latent Variables).
Let $P(X,S,Z)$ and $Q(X,S,Z)$ be two joint distributions over the observed variables $X$, proxy variables $S$, and latent variables $Z$. Suppose that the marginal distributions over $X$ and $S$ are identical, i.e., $P(X,S)=Q(X,S)$, and that Assumptions 1 through 4 are satisfied for both $P$ and $Q$. Then, there exists a permutation $\pi$ of $\{1,2,\dots,n_{z}\}$ such that for all $x\in\text{supp}(X)$ and for each $i\in\{1,2,\dots,n_{z}\}$, conditional distributions $P(Z\mid X=x)$ and $Q(Z\mid X=x)$ satisfy the following bound:
$\sup_{x\in\text{supp}(X)}\left|P(Z=i\mid X=x)-Q(Z=\pi(i)\mid X=x)\right|\leq O\left(\frac{1-\eta}{2\eta-1}\right)$ (1)
where $\eta\in\left(\frac{1}{2},1\right]$ is as defined in Assumption 4.
Proof Sketch
The proof proceeds in three stages as follows. 1) We first establish there is a unique value of $n_{z}$ that is compatible with a given observed distribution $P(X,S)$. Since $S$ is discrete, the conditional distribution $P(S\mid X=x)$ can be viewed as a vector for each $x\in\text{supp}(X)$. The collection of these vectors spans a subspace in $\mathbb{R}^{n_{s}}$. We show that the dimension of this subspace is $n_{z}$. 2) We show that the conditional distributions $P(Z\mid X=x)$ and $Q(Z\mid X=x)$ are related through a linear transformation. Since $S\perp X\mid Z$, we have: $P(S\mid X=x)=\sum_{z\in[n_{z}]}P(S\mid Z=z)P(Z=z\mid X=x)$ and similarly for $Q$. Because $P(S\mid X)=Q(S\mid X)$ for all $x$, it follows that there exists a matrix $A$ such that $P(Z\mid X=x)=A\!\ Q(Z\mid X=x)$. 3) Finally, using Assumption 4, which limits the overlap between latent variables, we show that $A$ can be approximated by a permutation matrix with an error bounded by $O\left(\frac{1-\eta}{2\eta-1}\right)$. This establishes a bound on the difference between $P(Z\mid X=x)$ and $Q(Z\mid X=x)$, yielding the approximate identifiability result. The formal proof is provided in Appendix A.1.
Full identifiability follows directly from Theorem 1.
Corollary 1 (Full Identifiability).
If $\eta=1$, then the latent variables are fully identifiable. That is, for all $x\in\text{supp}(X)$ and for each $i\in\{1,2,\dots,n_{z}\}$, there exists a permutation $\pi$ such that
$P(Z=i\mid X=x)=Q(Z=\pi(i)\mid X=x).$
Next, we argue that as the dimensionality of $X$ increases, $\eta\to 1$, making the latent distribution more identifiable. Intuitively, higher-dimensional covariates provide more information, making it easier to distinguish between the latent variable assignments, naturally driving $\eta$ closer to 1 and reducing overlap. The Weak Overlap assumption in Assumption 4 contrasts with the strict overlap assumption often used in causal inference literature, where it is typically assumed that $1-\eta\leq P(Z\mid X)\leq\eta$ for some $\eta\in(0,1)$ (van der Laan & Rose, 2011). It has been shown that as dimensionality increases, satisfying strict overlap becomes increasingly difficult (D’Amour et al., 2021). Inspired by this, we show that as the dimensionality of $X$ increases and the features in $X$ are sufficiently informative, $\eta=\max_{x}P(Z=j\mid X=x)$ approaches 1. This result builds on Proposition 2 of D’Amour et al. (2021), confirming that in high-dimensional spaces, the Weak Overlap assumption is easier to satisfy, leading to better identifiability of the latent variables.
Proposition 1.
Let $Z\in[n_{z}]$ and $X\in\mathbb{R}^{dim(X)}$. Further, assume that each feature $X_{i}$ is sufficiently discriminative of the latent values of $Z$, conditioned on all previous features $X_{1:i-1}$. Specifically, for all $i,j$, there exists $\epsilon>0$ such that, $\mathbb{E}_{P(X|Z=i)}\text{KL}\left(P(X_{i}|X_{1:i-1},Z=i)\parallel P(X_{i}|X_{1:i-1},Z=j)\right)$ is greater than $\epsilon$. Then, as $dim(X)\rightarrow\infty$, there exists no $\eta\in(0,1)$ such that
$\max_{x}P(Z=i\mid X=x)\leq\eta\quad\forall i,x.$
Additional Discussion of Assumptions
Assumptions 13 are common in the literature (Miao et al., 2018; Alabdulmohsin et al., 2023; Tsai et al., 2024). In contrast to prior works such as Tsai et al. (2024) and Miao et al. (2018), which depend on additional assumptions like the informative variable assumption (requiring that $P(Z\mid X)$ does not take a finite set of values) and the availability of both concept and proxy variables, our approach hinges on Assumption 4 for identifiability. We provide further discussion in Appendix B.
5 Methodology
In this section, we provide a formal description of our proposed framework, which involves two main phases: training and inference. Training Phase: We begin by estimating the latent distribution $P^{\text{tr}}(Z\mid X)$. To ensure identifiability of the latent distribution, we introduce a novel regularizer. Subsequently, we train a mixture-of-experts model to predict $Y$ from $X$ , where each expert corresponds to a specific latent confounder assignment. Inference Phase: In this phase, we estimate $P^{\text{te}}(Z)$, the marginal distribution of the confounder in the test data. Using this estimate, we obtain the shifted distribution $P^{\text{te}}(Y\mid X)$ by re-weighting the experts. This re-weighting adjusts the learned model to account for the shift in the latent confounder. These steps are illustrated in Figure 2.
Figure 2: This figure illustrates the methodology of our proposed model, which is structured into two primary phases: the Training Phase, where we first estimate the latent distribution and subsequently train a mixture-of-experts model for prediction; and the Testing Phase, where the model estimates $\frac{P^{\text{te}}(Z)}{P^{\text{tr}}(Z)}$ ($\mathbf{w}$) and reweights the gating function to predict the output for test input $X^{te}$.
5.1 Training
In training, we estimate the latent distribution $P^{\text{tr}}(Z\mid X)$ and train a mixture-of-experts predictor to predict the labels $Y$. The algorithm is the same for both proxy OOD and multi-source OOD tasks. We detail the process for the proxy OOD setting here, noting that the same steps apply to the multi-source OOD setting.
Estimating Latent Distribution:
Encoder-decoder models are often used to obtain the latent distributions (Kingma & Welling, 2013). We design a encoder-decoder model for our task leveraging the assumptions outlined in Section 4.
First, Assumption 1 $(S\perp X\mid Z)$ implies that $P^{\text{tr}}(S\mid X=x)$ can be factorized as follows:
$P^{\text{tr}}(S\mid X=x)=\sum_{z\in[n_{z}]}P(S\mid Z=z)P^{\text{tr}}(Z=z\mid X=x),$ (2)
This factorization allows us to decompose the distribution $P^{\text{tr}}(S\mid X)$ into two components: $P(S\mid Z)$ and $P^{\text{tr}}(Z\mid X)$. The latent variable $Z$ serves as an intermediate representation that captures the relevant information from $X$ about $S$. To accurately reconstruct $S$ from $X$, we need to model both $P^{\text{tr}}(Z\mid X)$ and $P(S\mid Z)$ effectively. We parameterize $P^{\text{tr}}(Z\mid X)$ using a neural network encoder $\phi:\mathbb{R}^{d_{x}}\rightarrow[0,1]^{d_{z}}$. The encoder transforms $X$ into a distribution over $Z$ using a softmax function in the final layer. The conditional distribution $P(S\mid Z)$, which is a discrete distribution, is then modeled by a matrix $M\in[0,1]^{d_{z}\times d_{s}}$, which acts as a decoder mapping $Z$ to $S$.
The parameters $\phi$ and $M$ can be learned by solving the following optimization problem:
$\phi^{*},M^{*}=\arg\min_{\phi,M}\mathbb{E}_{(X,S)\sim\mathcal{D}^{\text{tr}}}\left[\mathcal{L}_{S}(M^{\intercal}\phi(X),S)\right]$ (3)
where $M^{\intercal}\phi(x)=\sum_{z}\phi_{z}(x)M_{z,:}$, $\phi_{z}(x)$ denotes the $z^{\text{th}}$ component of $\phi(x)$, and $M_{z,:}$ represents the $z^{\text{th}}$ row of $M$ and $\mathcal{L}_{S}$ is the loss function.
Although the factorization in Equation 2 is not unique, meaning that various $M$ and $\phi$ can minimize the loss in Equation 3, enforcing Assumption 4 (weak overlap) for $\eta=1$ can help us uniquely identify $M$ and $\phi$. Recall that, for $\eta=1$, Assumption 4 implies for each confounder variable assignment, there exists $x\in\text{supp}(X)$ such that $P(Z=j\mid X=x)=1$. However, directly enforcing this assumption is infeasible due to the vast number of potential $X$ values to check for large datasets. Therefore, we design a regularization loss on the matrix $M$ which makes training tractable and gives us a condition that effectively enforces Assumption 4 to find $Z$ by regularizing the reconstruction loss in Equation 3.
Theorem 2.
Consider the set $\mathcal{M}$, which denotes the set of all matrices $M$ associated with some distribution $P(S\mid Z)$ for which there exists a $\phi(.)$ associated with density $P(Z\mid X)$ such that $M$ and $\phi(.)$ are minimizers of the reconstruction loss in Equation 3. Let $M^{*}\in\mathcal{M}$ be the matrix associated with a $Z$ that satisfies Assumption 4. Then it holds that
$\mathcal{L}_{\text{var}}(M^{*})\leq\mathcal{L}_{\text{var}}(M),\quad\forall M\in\mathcal{M},$
where
$\mathcal{L}_{\text{var}}(M)=\max_{z\in\{1,\dots,n_{z}\}}\left(\frac{1}{n_{s}}\sum_{s=1}^{n_{s}}\left(M_{s,z}-\frac{1}{n_{z}}\right)^{2}\right).$ (4)
In 2, each row $M_{i}$ represents $P(S\mid Z=i)$, and $\mathcal{L}_{\text{var}}(M)$ is the maximum variance among these rows. The result indicates that among all matrices $M$ that satisfy the factorization $M^{\top}\phi(x)=P(S\mid X=x)$ and minimize the reconstruction loss, the matrix $M^{*}$ that satisfies Assumption 4 will have the minimum highest variance of its rows. The formal proof is provided in Appendix A.4.
Therefore, our final optimization objective is
$\phi^{*},M^{*}=\arg\min_{\phi,M}L_{S}(M^{T}\phi(X),S)+\lambda\cdot L_{\text{var}}(M)$ (5)
with $\mathcal{L}_{S}$ as in Equation 3 and $\mathcal{L}_{\text{var}}$ as in Equation 13.
Mixture-of-Experts model.
In this phase, we use the estimated latent distribution $P^{\text{tr}}(Z\mid X)$ from the previous phase as a mixture distribution for our mixture-of-experts (MoE) model, which captures the heterogeneity of the prediction task.
MoE models consist of multiple expert neural networks $E_{k}(x)$ with a gating function to weigh and combine their outputs based on the input features (Jacobs et al., 1991). For our problem, each expert is specialized for a different $Z$, making $P(Y\mid X)$ different for each $Z$. The mixture distribution $P^{\text{tr}}(Z\mid X)$ determines the weights for combining the experts’ outputs. However, learning one distinct expert per value of $Z$ may lead to overfitting, especially for values of $Z$ that are associated with a small number of data samples. To mitigate this, we share all parameters except for the final layer of the experts. The shared layers learn a common representation of the input features $X$, following the intuition in Kirichenko et al. (2022) that lower layers can learn useful representations across different confounder assignments. The prediction for a particular input $x$ generated by our MoE model is given by:
$\text{MoE}(x)=\sum_{k=1}^{n_{z}}P(Z=k\mid X=x)E_{k}(h(x)),$
where $n_{z}$ is the number of experts. Here, $h(x)$ denotes the shared lower layers that learn a common representation of the input features, while $E_{k}(h(x))$ represents the output of the $k$-th expert network. The mixture weights $P(Z=k\mid X=x)$ determine the contribution of each expert’s output to the final prediction. Figure 2 shows an illustration of the architecture.
Training this mixture is key to our approach since it allows the model to capture heterogeneity more effectively than using $Z$ as a feature. This is because the mixture decomposes the prediction into the mixture $P(Z=k\mid X=x)$ and $P_{Z}(Y\mid X)$. In both of our tasks, a shift in the confounder only affects the gating function $P(Z=k\mid X=x)$, while $P_{Z}(Y\mid X)$ remains stable in both train and test. By adjusting and re-weighting the gating function, the model can respond to changes in the latent confounder.
5.2 Inference
During inference, there is a shift in $P(Z)$ from train to test, meaning that $P^{\text{tr}}(Z\mid X)\neq P^{\text{te}}(Z\mid X)$. This shift affects the mixture for our MoE model. Since $Z$ causes $X$, this scenario is analogous to a label shift task, where $P(X\mid Y)$ stays invariant, i.e., $P^{\text{tr}}(X\mid Z)=P^{\text{te}}(X\mid Z)$, but there is a shift in $P(Z)$ from train to test, i.e., $P^{\text{tr}}(Z)\neq P^{\text{te}}(Z)$. In this case, $P^{\text{te}}(Z\mid X)$ can be obtained by re-weighting $P^{\text{tr}}(Z\mid X)$ as follows:
$P^{\text{te}}(Z=z\mid X)$ $=\frac{P^{\text{tr}}(Z=z\mid X)\frac{P^{\text{te}}(Z=z)}{P^{\text{tr}}(Z=z)}}{\sum_{z^{\prime}}P^{\text{tr}}(Z=z^{\prime}\mid X)\frac{P^{\text{te}}(Z=z^{\prime})}{P^{\text{tr}}(Z=z^{\prime})}}$ (6)
This re-weighting, however, relies on the density ratio $\frac{P^{\text{te}}(Z)}{P^{\text{tr}}(Z)}$ ($\mathbf{w}$), which is directly unobserved but can be estimated from data. For instance, the method-of-moments approach introduced by Lipton et al. (2018) can be used for this estimation. The intuition behind their approach is that, since $Z$ causes $X$, $P(f(X)\mid Z)$ remains invariant from training to testing for any function $f(X)$. By observing the changes in the marginal distribution of $f(X)$, we can accurately estimate the shift in $P(Z)$.
The motivation for learning $P^{\text{te}}(Z\mid X)$ through re-weighting $P^{\text{tr}}(Z\mid X)$ instead of estimating $Z$ with our encoder-decoder framework from test data is twofold: 1) We aim to capture and adapt to changes in the underlying distribution of the latent confounder between training and test phases. 2) We do not assume access to proxy variables for proxy tasks and multi-source data during the test phase.
6 Experiments
In this section, we demonstrate the effectiveness of our approach on both synthetic and real data.
Baselines
We compare our method against several baselines, focusing on out-of-distribution (OOD) and domain adaptation techniques. For OOD methods, we include Invariant Risk Minimization (IRM) (Arjovsky et al., 2019), Variance Risk Extrapolation (VREx) (Krueger et al., 2021), and GroupDRO (Sagawa et al., 2020), which have access to the latent confounder during training. For domain adaptation, we consider Domain-Adversarial Neural Networks (DANN) (Ganin et al., 2016) and DeepCORAL (Sun & Saenko, 2016). Additionally, we include Proxy Domain Adaptation (ProxyDA) (Tsai et al., 2024) as a proxy-based domain adaptation method. It is important to note that some methods have access to test data during training: ProxyDA, DeepCORAL, and DANN use test data for domain adaptation purposes. Furthermore, GroupDRO, IRM, and VREx utilize ground-truth $Z$ labels during training, providing them with additional information about the dataset structure. ProxyDA requires both proxies and source/domain labels, but our approach only uses one. To ensure a fair comparison for tasks, we provide ProxyDA with random noise for the variable which is not available. Further details on the baselines and datasets can be found in Appendix C.
Synthetic Data
[…]
Performance on Synthetic Data
Table 1: Performance on Synthetic Data
Method Proxy OOD Task Multi-source OOD Task
ACC OOD $\uparrow$ ACC OOD $\uparrow$
ERM 0.484(0.001) 0.466 (0.006)
GroupDRO 0.465 (0.11) 0.443 (0.016)
IRM 0.465 (0.001) 0.453 (0.005)
DeepCORAL 0.463 (0.010) 0.464 (0.005)
VREx 0.523 (0.034) 0.490 (0.027)
DANN 0.467 (0.007) 0.459 (0.005)
ProxyDA 0.485 (0.047) 0.462 (0.012)
Ours ($\lambda=0$) 0.870 (0.021) 0.794 (0.037)
Ours 0.896 (0.002) 0.811 (0.039)
Table 1 shows the comparative performance on synthetic datasets. We report test accuracy on out-of-distribution tasks, presenting the mean accuracy over five runs along with the standard deviation. Our method significantly outperforms all baselines on both datasets. This is because most baselines rely on a single predictor, which cannot capture the heterogeneity in $P(Y\mid X)$. For ProxyDA, since only one of the proxy or multiple sources is available, it is unable to effectively adapt to the shift. Additionally, we observe an improvement in performance with the regularizer.
Real Data
For real data, we use datasets from Ding et al. (Ding et al., 2021), covering prediction tasks related to income, employment, health, transportation, and housing. These datasets span multiple years and all states of the United States, allowing researchers to study temporal shifts and geographic variations (Liu et al., 2024a). The input $X$ includes features such as education, occupation, marital status, and demographic attributes like sex and race.
For the Proxy OOD task, we evaluate using the ACS Employment dataset, which predicts employment status using 54 features. We consider disability status ($Z$) as the latent confounder. Here, $Z=1$ represents disabled individuals and $Z=0$ represents others. Proxies such as public insurance coverage and independent living status ($S$) are used to infer disability. The distribution of $Z$ during training is $P^{\text{tr}}(Z=0)=0.95$ and $P^{\text{tr}}(Z=1)=0.05$, while the test distribution flips to $P^{\text{te}}(Z=0)=0.05$ and $P^{\text{te}}(Z=1)=0.95$.
For the Multi-source OOD task, we use the ACS Income dataset to predict individual income using 76 features. We consider the state of residence ($Z$) as the latent confounder. We train on three synthetic splits: 20% Puerto Rico (PR) and 80% California (CA); 20% South Dakota (SD) and 80% California (CA); and 20% PR, 20% SD, and 60% CA. Testing is performed on a split of 90% PR, 5% SD, and 5% CA.
Performance on Real Data
Table 2: Performance on Real Data (Proxy and Multi-source OOD tasks)
Method ACS Employment Task ACS Income Task
ACC OOD $\uparrow$ ACC OOD $\uparrow$
ERM 0.670 (0.005) 0.805 (0.008)
GroupDRO 0.654 (0.014) 0.807 (0.026)
IRM 0.663 (0.006) 0.819 (0.022)
DeepCORAL 0.664 (0.007) 0.776 (0.021)
VREx 0.702 (0.014) 0.851 (0.008)
DANN 0.663 (0.007) 0.773 (0.031)
ProxyDA 0.689 (0.006) 0.856 (0.040)
Ours ($\lambda=0$) 0.710 (0.008) 0.856 (0.052)
Ours 0.709 (0.005) 0.883 (0.001)
Table 2 shows the comparative performance on ACS Employment and ACS Income. We report test accuracy on out-of-distribution tasks, presenting the mean accuracy over five runs along with the standard deviation. Notably, our method outperforms all baselines in both datasets, demonstrating its robust ability to handle shifts in $P(X,Y)$ caused by latent confounder shifts.
Scalability
Figure 3: Training time and memory usage for Our method and ProxyDA
To evaluate the scalability of our approach, we measured memory consumption and runtime during training on the Proxy task with synthetic data for different sample sizes. Figure 3 compares our method with ProxyDA, showing peak memory usage and runtime as functions of sample size. ProxyDA, as expected, shows cubic time complexity, typical of kernel methods. In contrast, our approach theoretically has linear time complexity for a fixed number of iterations. In practice, early stopping reduces the number of iterations for larger sample sizes, thereby lowering runtime further as seen in the plot. Regarding memory our method’s memory usage stays constant while ProxyDA’s memory usage increases with sample size. This makes it infeasible for ProxyDA to scale to datasets with large sample sizes.
7 Conclusion
In this work, we tackled the challenging task of out-of-distribution generalization in the presence of unobserved confounders. We introduced a simple yet scalable method that first estimates the distribution of latent confounder and then adapts the classifier to the test distribution. Our approach not only theoretically proves the identifiability of the latent distribution but also empirically outperforms existing baselines across multiple tasks. This demonstrates the robustness and practical effectiveness of our method, providing a strong foundation for future research in developing OOD-robust techniques that handle unobserved confounders. Our method operates under relatively simple assumptions, however, there may be scenarios where these assumptions fail to hold. Future work should consider understanding and identifying these limitations and build upon our proposed method.
Appendix A Proofs
A.1 Proof of Theorem 1
Theorem 1 (Approximate Identifiability of Latent Variables).
Let $P(X,S,Z)$ and $Q(X,S,Z)$ be two joint distributions over the observed variables $X$, proxy variables $S$, and latent variables $Z$. Suppose that the marginal distributions over $X$ and $S$ are identical, i.e., $P(X,S)=Q(X,S)$, and that Assumptions 1 through 3 are satisfied for both $P$ and $Q$. Then, there exists a permutation $\pi$ of $\{1,2,\dots,n_{z}\}$ such that for all $x\in\text{supp}(X)$ and for each $i\in\{1,2,\dots,n_{z}\}$, the conditional distributions $P(Z=i\mid X=x)$ and $Q(Z=\pi(i)\mid X=x)$ satisfy the following bound:
$\sup_{x\in\text{supp}(X)}\left|P(Z=i\mid X=x)-Q(Z=\pi(i)\mid X=x)\right|\leq O\left(\frac{1-\eta}{2\eta-1}\right)$ (7)
where $\eta\in\left(\frac{1}{2},1\right]$ is as defined in Assumption 4.
Proof.
The proof follows three main steps:
  1. 1.
    We first establish there is a unique value of $n_{z}$ that is compatible with a given observed distribution $P(X,S)$.
  2. 2.
    We show that the conditional distributions $P(Z\mid X=x)$ and $Q(Z\mid X=x)$ are related through a linear transformation. We further show that this linear transformation must satisfy certain constraints.
  3. 3.
    Finally, using Assumption 4, which limits the overlap between latent variables, we show that $A$ can be approximated by a permutation matrix with an error bounded by $O\left(\frac{1-\eta}{2\eta-1}\right)$. This establishes a bound on the difference between $P(Z\mid X=x)$ and $Q(Z\mid X=x)$.
1. Establishing Dimension of the Subspace Spanned by $P(S\mid X=x)$ = $n_{z}$
Under Assumption 1, we have that $S\perp X\mid Z$. Therefore, the conditional distribution can be expressed as:
$P(S\mid X=x)=\sum_{z=1}^{n_{z}}P(S\mid Z=z)P(Z=z\mid X=x).$
Let us consider $P(S\mid Z)$ as a matrix $M_{P}\in\mathbb{R}^{n_{z}\times n_{s}}$ where each row corresponds to $P(S\mid Z=z)$ for $z\in\{1,2,\dots,n_{z}\}$. By Assumption 2, the matrix $M_{P}$ has full rank, i.e., $\text{rank}(M_{P})=n_{z}$.
Since $P(S\mid X=x)$ is a linear combination of the rows of $M_{P}$, the set of all such vectors $\{P(S\mid X=x)\}_{x\in\text{supp}(X)}$ lies within the row space of $M_{P}$.
Given that $M_{P}$ has full rank, the dimension of this subspace is exactly $n_{z}$. Formally, the span of $\{P(S\mid X=x)\}$ satisfies:
$\text{dim}\left(\text{span}\left\{P(S\mid X=x)\mid x\in\text{supp}(X)\right\}\right)=n_{z}.$
Therefore, there is a unique $n_{z}$ for a given observed distribution.
2. Relating $P(Z\mid X=x)$ and $Q(Z\mid X=x)$ through a Linear Transformation
Given that $P(X,S)=Q(X,S)$, it follows that for all $x\in\text{supp}(X)$:
$P(S\mid X=x)=Q(S\mid X=x).$
Using Assumption 1, we can decompose these conditional distributions as:
$P(S\mid X=x)$ $=\sum_{z=1}^{n_{z}}P(S\mid Z=z)P(Z=z\mid X=x)=M_{P}^{T}P(Z\mid X=x),$ (8)
$Q(S\mid X=x)$ $=\sum_{z=1}^{n_{z}}Q(S\mid Z=z)Q(Z=z\mid X=x)=M_{Q}^{T}Q(Z\mid X=x),$ (9)
where we define the matrices $M_{P}\in\mathbb{R}^{n_{z}\times n_{s}}$ and $M_{Q}\in\mathbb{R}^{n_{z}\times n_{s}}$ such that:
$(M_{P})_{z,s}=P(S=s\mid Z=z),\quad(M_{Q})_{z,s}=Q(S=s\mid Z=z).$
Given that $P(S\mid X=x)=Q(S\mid X=x)$ for all $x$, we have:
$M_{P}^{T}P(Z\mid X=x)=M_{Q}^{T}Q(Z\mid X=x).$
Since Assumption 2 ensures that $M_{P}$ has full rank and Assumption 3 guarantees that $n_{s}\geq n_{z}$, the matrix $M_{P}^{\top}$ admits a left Moore-Penrose pseudo-inverse, denoted by $M_{P}^{\dagger}\in\mathbb{R}^{n_{z}\times n_{s}}$, satisfying:
$M_{P}^{\dagger}M_{P}^{\top}=I_{n_{z}},$
where $I_{n_{z}}$ is the $n_{z}\times n_{z}$ identity matrix. Multiplying both sides of the equation by $M_{P}^{\dagger}$ yields:
$M_{P}^{\dagger}M_{P}^{\top}P(Z\mid X=x)=M_{P}^{\dagger}M_{Q}^{\top}Q(Z\mid X=x),$
which simplifies to:
$P(Z\mid X=x)=AQ(Z\mid X=x),$ (10)
where we define the transformation matrix $A$ as:
$A=M_{P}^{\dagger}M_{Q}^{\top}.$
Constraints: We show that $A$ must satisfy certain constraints and cannot be an arbitrary matrix.
1)We claim that $\mathbf{1}^{\top}A=\mathbf{1}^{\top}$, where $\mathbf{1}\in\mathbb{R}^{n_{z}}$ is the vector of ones. In other words, the sum of elements in each column of $A$ is 1.
Proof: Given that $P(S\mid X=x)=Q(S\mid X=x)$ for all $x$, the columns of $M_{P}^{\top}$ and $M_{Q}^{\top}$ span the same subspace. Therefore, there exists a matrix $C\in\mathbb{R}^{n_{z}\times n_{z}}$ such that:
$M_{P}^{\top}C=M_{Q}^{\top}.$
This implies,
$C=M_{P}^{\dagger}M_{Q}^{\top}=A$
We know that both $M_{P}$ and $M_{Q}$ represent probability distributions $P(S\mid Z)$ and $Q(S\mid Z)$, meaning that:
$\mathbf{1}^{\top}M_{P}^{\top}=\mathbf{1}^{\top},\quad\mathbf{1}^{\top}M_{Q}^{\top}=\mathbf{1}^{\top},$
where $\mathbf{1}\in\mathbb{R}^{n_{s}}$ is the vector of ones. Applying these equalities to the equation $M_{P}^{\top}A=M_{Q}^{\top}$, we get:
$\mathbf{1}^{\top}M_{P}^{\top}A=\mathbf{1}^{\top}M_{Q}^{\top}.$
Substituting the known equalities:
$\mathbf{1}^{\top}A=\mathbf{1}^{\top},$ (11)
This means that the sum of the elements in each column of $A$ is 1.
2)We derive upper and lower bounds on the entries of $A$.
Consider any row $r$ of $A$. By the definition of conditional probabilities and equation equation 10, we have:
$P(Z=r\mid X=x)=\sum_{i=1}^{n_{z}}A_{ri}Q(Z=i\mid X=x),\quad\forall x\in\text{supp}(X).$
Since $P(Z=r\mid X=x)$ is a valid probability, it satisfies:
$0\leq\sum_{i=1}^{n_{z}}A_{ri}Q(Z=i\mid X=x)\leq 1$
Subtracting $\frac{1}{2}$ from both sides, we obtain:
$-\frac{1}{2}\leq\sum_{i=1}^{n_{z}}\left(A_{ri}-\frac{1}{2}\right)Q(Z=i\mid X=x)\leq\frac{1}{2}.$
Let $\lambda=\max_{i}\left|A_{ri}-\frac{1}{2}\right|$. Without loss of generality, assume that the maximum is achieved at $i=1$. Then:
$\left|\sum_{i=1}^{n_{z}}\left(A_{ri}-\frac{1}{2}\right)Q(Z=i\mid X=x)\right|\leq\frac{1}{2}.$
Applying the triangle inequality, we have:
$\left|\left(A_{r1}-\frac{1}{2}\right)Q(Z=1\mid X=x)\right|-\left|\sum_{i=2}^{n_{z}}\left(A_{ri}-\frac{1}{2}\right)Q(Z=i\mid X=x)\right|\leq\frac{1}{2}.$
Since $|A_{ri}-\frac{1}{2}|\leq\lambda$ for all $i$, it follows that:
$\lambda Q(Z=1\mid X=x)-\lambda\sum_{i=2}^{n_{z}}Q(Z=i\mid X=x)\leq\frac{1}{2}.$
Simplifying, we get:
$\lambda Q(Z=1\mid X=x)-\lambda(1-Q(Z=1\mid X=x))\leq\frac{1}{2},$
which further simplifies to:
$\lambda(2Q(Z=1\mid X=x)-1)\leq\frac{1}{2}.$
Given that there exists $x^{*}\in\text{supp}(X)$ such that $Q(Z=1\mid X=x^{*})\geq\eta$, where $\eta\in\left(\frac{1}{2},1\right]$, we substitute to obtain:
$\lambda(2\eta-1)\leq\frac{1}{2}\quad\Rightarrow\quad\lambda\leq\frac{1}{2(2\eta-1)}.$
Therefore, each entry $A_{ri}$ satisfies:
$-\frac{1-\eta}{2\eta-1}\leq A_{ri}\leq 1+\frac{1-\eta}{2\eta-1}.$ (12)
3)We derive lower bounds on the maximum entry in each row of $A$
Let $\mu=\max_{i}A_{ri}$. From the definition of $P(Z=r\mid X=x)$, we have:
$P(Z=r\mid X=x)=\sum_{i=1}^{n_{z}}A_{ri}Q(Z=i\mid X=x)\leq\mu\sum_{i=1}^{n_{z}}Q(Z=i\mid X=x)=\mu.$
Since $\max_{x}P(Z=r\mid X=x)\geq\eta$, it follows that $\mu\geq\eta$.
3. $A$ can be approximated by a permutation matrix
We aim to show that there exists a permutation matrix $\Pi$ such that:
$|A_{ij}-\Pi_{ij}|<O\left(\frac{1-\eta}{2\eta-1}\right),\quad\forall i,j.$
If $n_{z}=1$, A is trivially the identity matrix and is equal to a permutation matrix. Henceforth, we assume $n_{z}\geq 2$. Consider the following cases:
Case 1: $\frac{1}{2}<\eta\leq\frac{n_{z}+1}{n_{z}+2}$ For this case, consider any generic permutation matrix $\Pi$. From equation equation 12, we have:
$|A_{ij}-\Pi_{ij}|\leq 1+\frac{1-\eta}{2\eta-1}\leq\frac{n_{z}+2}{n_{z}+2}+\frac{1-\eta}{2\eta-1}\leq(n_{z}+2)(1-\eta)+\frac{1-\eta}{2\eta-1}\leq\frac{(n_{z}+2)(1-\eta)}{2\eta-1}+\frac{1-\eta}{2\eta-1}=O\left(\frac{1-\eta}{2\eta-1}\right).$
Case 2: $\frac{n_{z}+1}{n_{z}+2}<\eta\leq 1$
Starting with the given inequality $\eta>\frac{n_{z}+1}{n_{z}+2}$, we have
$\eta(n_{z}+2)$ $>n_{z}+1$
$n_{z}\eta-n_{z}$ $>1-2\eta$
$n_{z}(1-\eta)$ $<2\eta-1$
$(2\eta-1)$ $>(n_{z}-2)\frac{(1-\eta)}{2\eta-1}$
$2\eta$ $>1+(n_{z}-2)\frac{(1-\eta)}{2\eta-1}$
The above inequality, equation equation 11 and equation 12 imply that no more than one element in any column can be greater than or equal to $\eta$.
Since each row must have at least one element $\geq\eta$ and $A$ is a square matrix, it follows that each column has exactly one element $\geq\eta$. Without loss of generality, consider that for column $c$, the first element is $\geq\eta$, i.e. $A_{1c}\geq\eta$.
Then,
$\sum_{i=2}^{n_{z}}A_{ic}=1-A_{1c}$
Since all $A_{ic}\geq-\frac{1-\eta}{2\eta-1}$, we have:
$\max_{i\neq 1}A_{ic}\leq(1-\eta)+(n_{z}-2)\frac{1-\eta}{2\eta-1}$
Also,
$|1-A_{1c}|\leq\max\left\{1-\eta,\frac{1-\eta}{2\eta-1}\right\}$
and
$|A_{ic}|\leq\max\left\{(1-\eta)+(n_{z}-2)\frac{1-\eta}{2\eta-1},\frac{1-\eta}{2\eta-1}\right\}$
Define the permutation matrix $\Pi$ such that:
$\Pi_{ij}=\begin{cases}1&\text{if }A_{ij}\geq\eta,\\ 0&\text{otherwise}.\end{cases}$
By construction, $\Pi$ satisfies:
$|\Pi_{ij}-A_{ij}|\leq\max\left\{1-\eta,\frac{(n_{z}-2)(1-\eta)}{2\eta-1},\frac{(1-\eta)}{2\eta-1}\right\}=O\left(\frac{1-\eta}{2\eta-1}\right).$
We now obtain the final bound on the difference of the posterior distributions.
From equation equation 10, we have $P(Z|X=x)=AQ(Z|X=x)$. Using this, we can derive:
$|P(Z=i|X=x)-Q(Z=\pi(i)|X=x)|$ $=\left|\sum_{j=1}^{n_{z}}A_{ij}Q(Z=j|X=x)-Q(Z=\pi(i)|X=x)\right|$
$=\left|\sum_{j=1}^{n_{z}}(A_{ij}-\Pi_{ij})Q(Z=j|X=x)\right|$
$\leq\sum_{j=1}^{n_{z}}|A_{ij}-\Pi_{ij}||Q(Z=j|X=x)|$
$\leq\max_{j}|A_{ij}-\Pi_{ij}|\sum_{j=1}^{n_{z}}|Q(Z=j|X=x)|$
$=\max_{j}|A_{ij}-\Pi_{ij}|$
where we used the triangle inequality and the fact that $\sum_{j=1}^{n_{z}}|Q(Z=j|X=x)|=1$ since $Q(Z|X=x)$ is a probability distribution.
From our analysis in step 6, we know that:
$\max_{j}|A_{ij}-\Pi_{ij}|\leq O\left(\frac{1-\eta}{2\eta-1}\right)$
Therefore, we can conclude:
$\sup_{x\in\text{supp}(X)}\left|P(Z=i\mid X=x)-Q(Z=\pi(i)\mid X=x)\right|\leq O\left(\frac{1-\eta}{2\eta-1}\right),$
for a permutation $\pi$ corresponding to matrix $\Pi$
which proves Theorem 1. ∎
A.2 Proof of Corollary 1
Corollary 1 (Full Identifiability).
If $\eta=1$, then the latent variables are fully identifiable. That is, for all $x\in\text{supp}(X)$ and for each $i\in\{1,2,\dots,n_{z}\}$, there exists a permutation $\pi$ such that
$P(Z=i\mid X=x)=Q(Z=\pi(i)\mid X=x).$
Proof.
The proof follows directly from Theorem 1. When $\eta=1$, the bound in equation equation 7 becomes:
$\sup_{x\in\text{supp}(X)}\left|P(Z=i\mid X=x)-Q(Z=\pi(i)\mid X=x)\right|$ $\leq O\left(\frac{1-\eta}{2\eta-1}\right)$
$=O\left(\frac{1-1}{2(1)-1}\right)$
$=O(0)$
$=0$
Therefore,
$P(Z=i\mid X=x)=Q(Z=\pi(i)\mid X=x)$
for all $x$ and $i$, proving that the latent variables are fully identifiable up to permutation when $\eta=1$. ∎
A.3 Proof of Proposition 1
This proposition and proof closely follows D’Amour et al. (2021).
Proposition 1.
Let $Z\in[n_{z}]$ and $X\in\mathbb{R}^{dim(X)}$. Further, assume that each feature $X_{i}$ is sufficiently discriminative of the latent values of $Z$, conditioned on all previous features $X_{1:i-1}$. Specifically, for all $i,j$, there exists $\epsilon>0$ such that, $\mathbb{E}_{P(X|Z=i)}\text{KL}\left(P(X_{i}|X_{1:i-1},Z=i)\parallel P(X_{i}|X_{1:i-1},Z=j)\right)$ is greater than $\epsilon$. Then, as $dim(X)\rightarrow\infty$, there exists no $\eta\in(0,1)$ such that
$\max_{x}P(Z=i\mid X=x)\leq\eta\quad\forall i,x.$
Proof.
For simplicity, we assume $Z\in\{0,1\}$. We proceed by contradiction.
Suppose there exists an $\eta\in(0,1)$ such that for all $x$:
$P(Z=0\mid X=x)\leq\eta\quad\text{and}\quad P(Z=1\mid X=x)\leq\eta$
This implies that for all x:
$1-\eta\leq P(Z=0\mid X=x)\leq\eta\quad\text{and}\quad 1-\eta\leq P(Z=1\mid X=x)\leq\eta$
It follows that $P(Z=0)>0$ and $P(Z=1)>0$. Let $\gamma=P(Z=1)/P(Z=0)$. For any $x$:
$\frac{P(X=x\mid Z=0)}{P(X=x\mid Z=1)}=\frac{P(Z=0\mid X=x)}{P(Z=1\mid X=x)}\cdot\frac{1}{\gamma}\leq\frac{\eta}{1-\eta}\cdot\frac{1}{\gamma}$
Taking logarithms:
$\log\frac{P(X=x\mid Z=0)}{P(X=x\mid Z=1)}\leq\log\frac{\eta}{1-\eta}-\log\gamma$
Now, consider the KL divergence between $P(X\mid Z=0)$ and $P(X\mid Z=1)$:
$\mathrm{KL}(P(X\mid Z=0)\parallel P(X\mid Z=1))$
$=\mathbb{E}_{X\sim P(X\mid Z=0)}\left[\log\frac{P(X\mid Z=0)}{P(X\mid Z=1)}\right]$
$\leq\log\frac{\eta}{1-\eta}-\log\gamma$
This implies that the KL divergence is bounded. However, by the chain rule of KL divergence and the condition given in the proposition:
$\mathrm{KL}(P(X\mid Z=0)\parallel P(X\mid Z=1))$
$=\sum_{i=1}^{\dim(X)}\mathbb{E}_{P(X\mid Z=0)}\left[\mathrm{KL}(P(X_{i}\mid X_{1:i-1},Z=0)\parallel P(X_{i}\mid X_{1:i-1},Z=1))\right]$
$\geq\dim(X)\cdot\epsilon$
As $\dim(X)\to\infty$, the right-hand side approaches infinity, contradicting the boundedness we established earlier. Therefore, our initial assumption must be false. We conclude that as $\dim(X)\to\infty$, there exists no $\eta\in(0,1)$ such that $\max_{x}P(Z=z\mid X=x)\leq\eta$ for all $z\in\{0,1\}$ ∎
A.4 Proof of Theorem 2
Theorem 2.
Consider the set $\mathcal{M}$, which denotes the set of all matrices $M$ associated with some distribution $P(S\mid Z)$ for which there exists a $\phi(.)$ associated with density $P(Z\mid X)$ such that $M$ and $\phi(.)$ are minimizers of the reconstruction loss $\mathbb{E}_{(X,S)\sim\mathcal{D}^{\text{tr}}}\left[\mathcal{L}_{S}(M^{\intercal}\phi(X),S)\right]$. Let $M^{*}\in\mathcal{M}$ be the matrix associated with a $Z$ that satisfies Assumption 4 for $\eta=1$. Then it holds that
$\mathcal{L}_{\text{var}}(M^{*})\leq\mathcal{L}_{\text{var}}(M),\quad\forall M\in\mathcal{M},$
where
$\mathcal{L}_{\text{var}}(M)=\max_{z\in\{1,\dots,n_{z}\}}\left(\frac{1}{n_{s}}\sum_{s=1}^{n_{s}}\left(M_{s,z}-\frac{1}{n_{s}}\right)^{2}\right).$ (13)
Proof.
To prove this, we first prove the following lemma.
Lemma 1.
Let $\mathbf{v},\mathbf{w}_{1},\mathbf{w}_{2},\ldots,\mathbf{w}_{n}\in[0,1]^{k}$ be vectors representing probability distributions, i.e., $\sum_{j=1}^{k}w_{ij}=1$ for all $1\leq i\leq n$ and $\sum_{j=1}^{k}v_{j}=1$. Suppose $\mathbf{v}=\sum_{i=1}^{n}\alpha_{i}\mathbf{w}_{i}$ where $\sum_{i=1}^{n}\alpha_{i}=1$ and $\alpha_{i}\geq 0$ for all $i$. That is, $\mathbf{v}$ is a convex combination of the vectors $\mathbf{w}_{i}$. Let $\text{Var}(\mathbf{v})$ denote the variance of this categorical distribution. Under these conditions, the variance of $\mathbf{v}$ satisfies:
$\text{Var}(\mathbf{v})\leq\sum_{i=1}^{n}\alpha_{i}\text{Var}(\mathbf{w}_{i})\leq\max_{i}\text{Var}(\mathbf{w}_{i}).$
Proof.
Recall the definition of variance for a vector $\mathbf{v}\in\mathbb{R}^{k}$ is given by:
$\text{Var}(\mathbf{v})=\frac{1}{k}\sum_{j=1}^{k}(v_{j}-\bar{v})^{2},$
where $v_{j}$ is the $j$-th component of $\mathbf{v}$ and $\bar{v}$ is the mean of the components of $\mathbf{v}$. Given $\mathbf{v}=\sum_{i=1}^{n}\alpha_{i}\mathbf{w}_{i}$, each component $v_{j}$ of $\mathbf{v}$ can be expressed as:
$v_{j}=\sum_{i=1}^{n}\alpha_{i}w_{ij}.$
The mean $\bar{v}$ of the components of $\mathbf{v}$ can be calculated as:
$\bar{v}=\frac{1}{k}\sum_{j=1}^{k}v_{j}=\frac{1}{k}\sum_{j=1}^{k}\sum_{i=1}^{n}\alpha_{i}w_{ij}.$
Since each $\mathbf{w}_{i}$ sums to 1, it follows that:
$\bar{v}=\sum_{i=1}^{n}\alpha_{i}\left(\frac{1}{k}\sum_{j=1}^{k}w_{ij}\right)=\sum_{i=1}^{n}\alpha_{i}\bar{w_{i}}.$
Now, substituting back into the variance formula, we get:
$\text{Var}(\mathbf{v})=\frac{1}{k}\sum_{j=1}^{k}\left(\sum_{i=1}^{n}\alpha_{i}(w_{ij}-\bar{w_{i}})\right)^{2}.$
Applying the AM-RMS inequality:
$\left(\sum_{i=1}^{n}\alpha_{i}(w_{ij}-\bar{w_{i}})\right)^{2}\leq\left(\sum_{i=1}^{n}\alpha_{i}(w_{ij}-\bar{w_{i}})^{2}\right),$
and since $\alpha_{i}^{2}\leq\alpha_{i}\in[0,1]$, we can simplify the variance expression:
$\text{Var}(\mathbf{v})\leq\sum_{i=1}^{n}\alpha_{i}\frac{1}{k}\sum_{j=1}^{k}(w_{ij}-\bar{w_{i}})^{2}.$
Finally, as $\frac{1}{k}\sum_{j=1}^{k}(w_{ij}-\bar{w_{i}})^{2}$ is the expression for the variance of $\mathbf{w}_{i}$, we conclude:
$\text{Var}(\mathbf{v})\leq\sum_{i=1}^{n}\alpha_{i}\text{Var}(\mathbf{w}_{i})\leq\max_{i}\text{Var}(\mathbf{w}_{i}).$
Now, we discuss the proof of Theorem 2.
Since $(\phi^{*}(.),M^{*})$ satisfies the weak overlap Assumption 4 for $\eta=1$, we know that $\exists x^{*}$ such that $\phi^{*}_{i}(x^{*})=1$ for all $i,1\leq i\leq n_{z}$ (because $\phi(x)=P(Z\mid X=x)$). Therefore, we know that
$\forall i,\exists x^{*}\text{ such that }P(S\mid X=x^{*})=\sum_{z}\phi^{*}_{z}(x^{*})M^{*}_{z,:}=M^{*}_{i,:}\,.$
Moreover, for any valid $\phi_{z}(x^{*})$,
$P(S\mid X=x^{*})=\sum_{z}\phi_{z}(x^{*})M_{z,:}\,.$
Therefore, for all i,
$M^{*}_{i,:}=\sum_{z}\phi_{z}(x^{*})M_{z,:}\,.$
Hence, using Lemma 1, for all $i$ we have
$\text{Var}(M^{*}_{i,:})\leq\max_{z}\text{Var}(M_{z,:})\,.$
Hence,
$\max_{z}\text{Var}(M^{*}_{z,:})\leq\max_{z}\text{Var}(M_{z,:}).$
Since $M^{*}_{z,:}$ and $M_{z,:}\in\mathbb{R}^{n_{s}}$, the mean of these is $\frac{1}{n_{s}}$. Hence,
$\max_{z}\text{Var}(M_{z,:})=\mathcal{L}_{\text{var}}(M)=\max_{z\in\{1,\dots,n_{z}\}}\left(\frac{1}{n_{s}}\sum_{s=1}^{n_{s}}(M_{s,z}-\frac{1}{n_{s}})^{2}\right),$
$\max_{z}\text{Var}(M^{*}_{z,:})=\mathcal{L}_{\text{var}}(M^{*})=\max_{z\in\{1,\dots,n_{z}\}}\left(\frac{1}{n_{s}}\sum_{s=1}^{n_{s}}(M^{*}_{s,z}-\frac{1}{n_{s}})^{2}\right).$
Therefore,
$\mathcal{L}_{\text{var}}(M^{*})\leq\mathcal{L}_{\text{var}}(M).$
Appendix B Detailed discussion on assumptions
In this section, we comprehensively discuss our assumptions and compare them to those in our closest related work, Proxy Domain Adaptation (ProxyDA) (Tsai et al., 2024). We then examine the practicality of these assumptions. Note that ProxyDA operates under two settings: one with multiple sources and a proxy variable, and the other with a concept variable and a proxy variable. Since our model is designed to work with either multiple sources or a proxy variable, we will only compare it with the former setting.
Assumptions Ours Proxy Domain Adaptation
Task Out-of-distribution Domain Adaptation
Observed variables $X$, $Y$ and ($U$ or $S$) $X$, $Y$, $U$ and $S$
Number of labeled sources 1 multiple ($\geq n_{z}$)
Conditional independence $S\perp X,Y\mid Z$ (or $U\perp X,Y\mid Z$) $U\perp X,Y,S\mid Z$ and $S\perp X\mid Z$
Causality between $X$ and $Y$ $X\leftrightarrow Y$ $X\rightarrow Y$
Full rank $P(S\mid Z)$ (or $P(U\mid Z)$) must be full rank $P(S\mid Z)$ must be full rank and $P(Z\mid U,X)$ must be full rank for X
Discrete unobserved confounder Yes Yes
Informative variable No Yes
Existence of high-likelihood sample Yes No
Proxy available at test time No Yes
$ns\geq n_{z}$ Yes Yes
$n_{z}$ known No No
Table 3: Comparison of Assumptions between Ours and Tsai et al. (2024)
Our assumptions
(i) In real-world tasks, test data is rarely available during training and train data is rarely available during inference. Hence, the OOD task we consider is more realistic and challenging. (ii) We only require $U$ (multiple training sources) or $S$, a proxy for the confounder, not both simultaneously. Additionally, we do not require a proxy at test time, relaxing previous assumptions. Moreover, for the multi-source task, only one of the sources needs to be labeled. (iii) Unlike ProxyDA, we allow for a bi-directional arrow between $X$ and $Y$ in our causal diagram. The causal relationship between $X$ and $Y$ can vary widely based on the problem. For example, in a house price prediction problem, it is likely that $X\rightarrow Y$ since features of the house likely causally affect the price of the house. However, in image classification problems, it is common to have $Y\rightarrow X$, where the category of the object causes the distribution of pixels in the image. Furthermore, in many cases, some features in $X$ might cause $Y$ and others may be effects of $Y$. For example, in our income prediction task, it is likely that education causes income but health insurance coverage is caused by income. Both are features in $X$. Therefore, allowing a bidirectional causal relationship between $X$ and $Y$ increases the scope of our method. (iv) We are slightly more restrictive regarding conditional independence. While ProxyDA allows $S$ to cause $Y$, we do not allow that relation. If $S$ causes $Y$, we cannot allow $Y$ to cause $X$. This is because if $S$ causes $Y$ and $Y$ causes $X$, then $S\perp X\mid Z$ does not hold. Therefore, there is an inherent trade-off between (iii) and (iv).
(v)As far as we know, all latent shift methods (Alabdulmohsin et al., 2023; Tsai et al., 2024) make full rank assumptions. Full rank assumptions ensure diversity in distributions with respect to the confounder and are necessary to prove identifiability. However, we do not require any full rank condition on $P(Z\mid S,X)$ or $P(Z\mid X)$ for all $X$, which could be quite strict. For example, the full rank assumption on $P(Z\mid S,X)$ for all $X$ does not allow $Z$ to be deterministic for any $X$.
(vi)While the informative variable assumption is common in the literature (Miao et al., 2018; Tsai et al., 2024), we do not require this assumption. This assumption does not allow certain distributions, for example, those where $P(Z\mid X)$ can take only a finite set of values.
One of the key assumptions we make, which other papers do not, is the weak overlap assumption. Our approximate identifiability result in Theorem 1 shows that we can bound the difference between the posteriors of any two possible distributions by $O(\frac{1-\eta}{2\eta-1})$. Proposition 1 shows that Assumption 4 would hold for an $\eta$ arbitrarily close to 1 for high-dimensional data. Therefore, given the observed distribution $P(X,S)$, we can approximate the latent distribution by an arbitrarily small error with the increase in dimension of $X$.
Appendix C Experimental Details
In this section, we provide additional details on our dataset construction, baselines, model implementation, and model selection.
C.1 Synthetic Data
Proxy OOD Task
For the proxy OOD task, we define the latent confounder $Z\in\{0,1,2\}$ and the proxy $S\in\{0,1,2\}$ with $X\in\mathbb{R}^{3}$ as the observed feature vector. The conditional distribution of $S$ given $Z$, denoted as $P(S\mid Z)$, is modeled with different probabilities depending on the value of $Z$. Specifically, the conditional probabilities for each value of $Z$ are given by:
$P(S=0\mid Z=0)=0.7,\quad P(S=1\mid Z=0)=0.3,$
$P(S=0\mid Z=1)=0.8,\quad P(S=1\mid Z=1)=0.1,\quad P(S=2\mid Z=1)=0.1,$
$P(S=0\mid Z=2)=0.1,\quad P(S=2\mid Z=2)=0.9.$
The feature vector $X=(X_{1},X_{2},X_{3})$ is generated conditionally on $Z$ as follows:
$X_{1}=-Z_{x}+Z+\epsilon_{1},\quad X_{2}=2Z_{x}+\epsilon_{2},\quad X_{3}=3Z_{x}+\epsilon_{3},$
where $Z_{x}\sim\mathcal{N}(0,1)$ is a Gaussian random variable, and $\epsilon_{1}\sim\mathcal{N}(0,0.2)$, $\epsilon_{2}\sim\mathcal{N}(0,1)$, $\epsilon_{3}\sim\mathcal{N}(0,1)$ are noise terms.
Finally, the binary target variable $Y$ is generated as a Bernoulli random variable based on a logistic function of $X$ and $Z$. The conditional probability of $Y$ given $X$ and $Z$ is:
$P(Y=1\mid X,Z)=\sigma(f(X,Z)),$
where $\sigma(\cdot)$ is the sigmoid function and the argument $\text{prob}(X,Z)$ is:
$f(X,Z)=\begin{cases}-0.6X_{1}+2.7X_{2}+0.6X_{3}-0.6Z+1.5&\text{if }Z=0,\\ 1.5X_{1}+(1.2-6Z)X_{2}+2.1X_{3}-0.6Z+1.5&\text{if }Z\neq 0.\end{cases}$
For the training set, the latent variable $Z$ is sampled from a multinomial distribution with probabilities $P(Z)=[0.15,0.35,0.5]$. For the test set, the probabilities are shifted to $P(Z)=[0.7,0.2,0.1]$ to simulate distribution shift.
Multi-source OOD Task
For the multi-source OOD task, we also define $Z\in\{0,1,2\}$ as the latent confounder and $X\in\mathbb{R}^{3}$ as the feature vector. The conditional distribution $P(X\mid Z)$ follows the same structure as in the proxy OOD task.
The difference lies in how the training and test distributions for $Z$ are defined. During training, we use multiple source domains, each with different distributions of $Z$. For example, we sample from three source domains with distributions:
$P(Z)=[0,0.7,0.3],\quad P(Z)=[0.8,0.1,0.1],\quad P(Z)=[0.1,0.0,0.9].$
In the test domain, the distribution shifts to:
$P(Z)=[0.7,0.2,0.1].$
The target variable $Y$ is generated similarly to the proxy OOD task, with $P(Y\mid X,Z)$ following a logistic function based on $X$ and $Z$.
C.2 Real Datasets
We construct our datasets using the folktables Python package (Ding et al., 2021). Below, we provide detailed information about each dataset and the rationale behind our splits.
ACSEmployment
The ACSEmployment dataset is used to predict whether an individual is employed. Disability status is the latent confounder ($Z$), where $Z=1$ denotes individuals with disabilities, and $Z=0$ denotes individuals without disabilities. This setup models real-world conditions where disabled individuals might face different employment opportunities and challenges. Public insurance coverage and independent living status serve as proxies for disability status. In the training set, $P^{\text{tr}}(Z=0)=0.95$ and $P^{\text{tr}}(Z=1)=0.05$, while in the test set, $P^{\text{te}}(Z=0)=0.05$ and $P^{\text{te}}(Z=1)=0.95$. Both the training and test sets contain 170,000 samples each. The input has 54 features.
ACSIncome
The ACSIncome dataset is used to predict an individual’s income. The state of residence ($Z$) is the latent confounder. We created three synthetic splits for training:
  • 20% Puerto Rico (PR) and 80% California (CA)
  • 20% South Dakota (SD) and 80% California (CA)
  • 20% PR, 20% SD, and 60% CA
Testing is performed on a split of 90% PR, 5% SD, and 5% CA. The training set comprises 10,502 samples, while the test set contains 3,481 samples. The input has 76 features.
To create significant shifts in $P(Y\mid X)$ across subpopulations, we first trained a classifier on data from California and tested it on data from all other states. Additionally, we trained a classifier for each state and tested it within the same state. We selected Puerto Rico and South Dakota as they exhibited the largest accuracy disparities compared to California. This method simulates real-world scenarios where economic conditions and job markets vary significantly across states, thereby affecting income levels.
C.3 Implementation Details
The training was performed on an 11th Gen Intel(R) Core(TM) i7-11390H CPU and on an Intel(R) Xeon(R) Gold 6230 CPU @ 2.10GHz. The experiments took approximately 48 hours to run. Our architecture employs a two-hidden-layer neural network for $\phi(\cdot)$. Furthermore, we use a Batch Normalization layer before the softmax layer to stabilize training.
We impose a constraint on the matrix $M$ such that the sum of the values in each row equals 1, and we clip all values to remain within the range of 0 to 1. In Theorem 2, we show how encouraging distributions with low variance leads to recovering the ground-truth latent distribution. The regularization parameter $\lambda$ is tuned on the validation set from {1e-3,1e-2,1e-1,1e0,1e1}.
For optimization, we utilize the Adam optimizer with a learning rate of $1\times 10^{-3}$. While training $\phi(\cdot)$, we run the model for 100 epochs and perform early stopping based on a held-out validation set from the same distribution as the training set.
For the mixture of experts model, we train the model for 25 epochs with early stopping based on a validation set from the same distribution as the training set.
Baselines
We implement GroupDRO (Sagawa et al., 2020), IRM (Arjovsky et al., 2019), V-REx (Krueger et al., 2021), DeepCORAL (Sun & Saenko, 2016), and DANN (Ganin et al., 2016) using the DomainBed code (Gulrajani & Lopez-Paz, 2020). We use early stopping based on a validation set for all these baselines. We implement ProxyDA  (Tsai et al., 2024) using their publicly available GitHub repository. For V-REx, IRM, and DeepCORAL, we tune the regularization penalty from {1e-2, 1e-1, 1e0, 1e1}. For GroupDRO, we tune the group-step size from {1e-2, 1e-1, 1e0, 1e1}. For ProxyDA, we used random noise for the other variable. We subsampled all training datasets on the real tasks to 5000 samples respectively due to memory constraints. We tuned three hyperparameters: cme, m0, and scale using a validation set from the same distribution as the training set, within the range {1e-4, 1e-3, 1e-2, 1e-1, 1e0}. For the ACS Employment dataset, we transformed the data into a lower dimension using Gaussian Random Projection as suggested by Tsai et al. (2024), tuning the number of dimensions from {2, 4, 8, 16}.